A submarine can stand a pressure of 1400 kN/m² above atmospheric pressure.

find the maximum depth at wich she can work in sea water of RD=1.025



1
Expert's answer
2021-06-14T08:51:36-0400

The maximum depth can be found from the maximum pressure by the following law:


"p_\text{max}=\rho gh,\\\space\\
h=\frac{p_\text{max}}{\rho g}=\frac{1400·10^3}{9.8·(1.025·1000)}=139\text{ m}."